OpenStudy (anonymous):

Given the function g(x) = −4x + 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values. a.The value of g(2) is larger than the value of g(−4). b. The value of g(2) is smaller than the value of g(−4). C. The value of g(2) is the same as the value of g(−4). d. The values of g(2) and g(−4) cannot be compared.

OpenStudy (anonymous):

g(x) = −4x + 5, compare and contrast g(2) and g(−4) g(2) = −4(2) + 5 g(2) = −8 + 5 = -3 g(-4) = −4(-4) + 5 g(-4) = 16 + 5 = 21

OpenStudy (anonymous):

b. The value of g(2) is smaller than the value of g(−4).

OpenStudy (anonymous):

thank you!